Suppose that the functions f and g are defined as follows
f(x)=(1)/(5x^(2)+4)
g(x)=(1)/(√(3x+5))
Find f*g and f+g . Then, give their domains using interval notation.
(f*g)(x)=
domain of f*g =
(f+g)(x)=
domain of f+g =
Any help would be appreciated.
Thank you

Answer by oscargut(2103)

Suppose that the functions f and g are defined as follows
f(x)=(1)/(5x^(2)+4)
g(x)=(1)/(√(3x+5))

Find f*g and f+g .

Then, give their domains using interval notation.
(f*g)(x)= 1/(5x^(2)+4)(√(3x+5))
domain of f*g = {x/ x > -5/3}
(f+g)(x)=(1)/(5x^(2)+4)+(1)/(√(3x+5))
domain of f+g = {x/ x > -5/3}
I can help you with more questions at: mthman@gmail.com
Thanks